OpenStudy (anonymous):

Can you guys help me please?? cos(pi/6 + (inverse)cos 3/4)

OpenStudy (anonymous):

\[\cos(x+y)=\cos(x)\cos(y)-\sin(x)\sin(y)\]

OpenStudy (anonymous):

\[\cos(\frac{\pi}{6}+\cos^{-1}(x))=\cos(\frac{\pi}{6})\cos(\cos^{-1}(x))-\sin(\frac{\pi}{6})\sin(\cos^{-1}(x))\]

OpenStudy (anonymous):

some of these are numbers you may know for example \(\cos(\frac{\pi}{6})=\frac{\sqrt3}{2}\)

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

also \[\cos^{-1}(\cos(x))=x\]

OpenStudy (anonymous):

\[\sin(\frac{\pi}{6})=\frac{1}{2}\]

OpenStudy (anonymous):

only one left to figure out is what is \[\sin(\cos^{-1}(x))\]

OpenStudy (anonymous):

you know what that is?

OpenStudy (anonymous):

oooh i misread the question the second part is \(\cos^{-1}(\frac{4}{5})\) right

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

even easier \[\cos(\frac{\pi}{6})\cos(\cos^{-1}(\frac{4}{5}))-\sin(\frac{\pi}{6})\sin(\cos^{-1}(\frac{4}{5}))\]

OpenStudy (anonymous):

I thought that you do \[\sin(\cos ^{-1}3/4)\]

OpenStudy (anonymous):

which = sin v

OpenStudy (anonymous):

yeah that is all that is left

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

we got \[\frac{\sqrt3}{2}\times \frac{3}{4}-\frac{1}{2}\times \sin(\cos^{-1}(\frac{3}{4})\]so far

OpenStudy (anonymous):

soo.. \[\sqrt{1-\cos^2}v ?\]

OpenStudy (anonymous):

yeah that will do it i do it by drawing a triangle

OpenStudy (anonymous):

ooo I wanna draw a triangle how do you do it that way?

OpenStudy (anonymous):



OpenStudy (anonymous):

you're so cool..

OpenStudy (anonymous):

there is a picture of an angle whose cosine is \(\frac{3}{4}\) all that is missing is the third side, which you find via pythagoras it is \[\sqrt{4^2-3^2}=\sqrt7\]

OpenStudy (anonymous):

yeah cool, why i am on dorky math site at night, but thanks...

OpenStudy (anonymous):

haha I think you're awesome :p

OpenStudy (anonymous):



OpenStudy (anonymous):

soo you get sqrt(7)/4?

OpenStudy (anonymous):

making \[\sin(\cos^{-1}(\frac{3}{4}))=\frac{\sqrt7}{4}\]

OpenStudy (anonymous):

right what you said

OpenStudy (anonymous):

B)

OpenStudy (anonymous):

you will get the same answer the other way

OpenStudy (anonymous):

the rest you basically plug in and its all algebra right?

OpenStudy (anonymous):

shush

OpenStudy (anonymous):

i would call it arithmetic, but yes, just plug them in

OpenStudy (anonymous):

i wanna do it :P

OpenStudy (anonymous):

fine good idea let me know what you get

OpenStudy (anonymous):

is it \[(3\sqrt{3}-\sqrt{7})/8\]?

OpenStudy (anonymous):

not telling

OpenStudy (anonymous):

comeonnnnn

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

ok yes it is

OpenStudy (anonymous):

it has to beee

OpenStudy (anonymous):

yay <3

OpenStudy (anonymous):

yes you are right

OpenStudy (anonymous):

haha thanks :)

OpenStudy (anonymous):

what's your name mister satellite?

OpenStudy (anonymous):

never seen some one excited over a trickonometry problem glad you enjoy them

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

" trickonometry"

OpenStudy (anonymous):

actually don't tell me your name you might get in trouble :P

OpenStudy (anonymous):

Bye :p

OpenStudy (anonymous):

i am just satellite73 bye

OpenStudy (anonymous):

You're just like a superhero... just satellite73 :p

OpenStudy (anonymous):

yikes how many picture you got?

OpenStudy (anonymous):

sorry haha :p

OpenStudy (anonymous):

just gotta find the right one ya know :p

OpenStudy (anonymous):

this looks good